Given n 4 n for a fixed positive integer n ≥ 2, prove that (n + 1) 4 n + 1.

Given n4 n for a fixed positive integer n ≥ 2, prove that (n + 1)4 n + 1.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. Given that n4 < 10n for n for a fixed + ve integer n ≥2.
To prove that (n + 1)4 , 10n + 1
Proof : Since n4 < 10n ⇒ 10n4 < 10n + 1 . . . . . . . . . . . . . . . . . . (1)
So it is sufficient to prove that (n + 1)4 < 10n4
Now (n + 1/n)4 = (1 + 1/n)4 ≤ (1 + 1/2)4 [∵ n ≥ 2]
= 81/16 < 10
⇒ (n + 1)4 < 10n4 . . . . . . . . . . . . . . . . . . . . . . . .(2)
From (1) and (2), (n + 1)4 < 10n + 1